On a map, two cities are 1 1/2 in. apart. If 3/8 in. on the map represents 45 mi, find the number of miles between the two cities.
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

.375/45=1.5/x
.375x=67.5
x=180 miles

Answer by stanbon(75887)

On a map, two cities are 1 1/2 in. apart. If 3/8 in. on the map represents 45 mi, find the number of miles between the two cities.
------------------------
miles/inches = miles/inches
----------------------
m/(3/2) = 45/(3/8)
----
m = (3/2)[15*8)
-----
m = 180 miles
==================
Cheers,
Stan H.